### v2

# Template-free and Rig-free Animation Transfer using Kinetic Codes

Abstract.
Animation retargetting applies sparse motion description (e.g., keypoint sequences) to a character mesh to produce a semantically plausible and temporally coherent full-body mesh sequence. Existing approaches come with restrictions – they require access to template-based shape priors or artist-designed deformation rigs, suffer from limited generalization to unseen motion and/or shapes, or exhibit motion jitter. We propose Self-supervised Motion Fields (SMF), a self-supervised framework that is trained with only sparse motion representations, without requiring dataset-specific annotations, templates, or rigs. At the heart of our method are Kinetic Codes, a novel autoencoder-based sparse motion encoding, that exposes a semantically rich latent space, simplifying large-scale training. Our architecture comprises dedicated spatial and temporal gradient predictors, which are jointly trained in an end-to-end fashion. The combined network, regularized by the Kinetic Codes’ latent space, has good generalization across both unseen shapes and new motions. We evaluated our method on unseen motion sampled from AMASS, D4D, Mixamo, and raw monocular video for animation transfer on various characters with varying shapes and topology. We report a new SoTA on the AMASS dataset in the context of generalization to unseen motion. Code, weights, and supplementary are available on the project webpage at https://motionfields.github.io/.
1. Introduction
Motion brings characters to life. To build any digital twin world, be it for scenario planning, games, or movies, adding motion to static characters is a fundamental requirement. Manually authoring full-body motion is tedious, error-prone, and requires significant effort from skilled artists. This quickly becomes expensive when scaling to long sequences or animating many different characters. Hence, researchers seek data-driven solutions.
Traditional approaches investigate this problem with explicit priors — statistical body templates (e.g., SMPL (Loper et al., 2015) for humans, CAFM (Sun and Murata, 2020) for animals) to model shape and pose variations or constructing specific character rigs to transfer joint motion to full body motion via skinning weights. These approaches are simple, popular, and robust, but come at the cost of having to first build an expressive shape space along with a corresponding pose space, and are restricted to specific templates.
Automating these workflows with learning-based solutions has gained popularity: learning a space of stick figure character motion (e.g., neural motion field (He et al., 2022)) or phase-based character control (Holden et al., 2017)); enabling deformation transfer from a source-target pair to a new shape via neural Jacobian fields (Aigerman et al., 2022; Muralikrishnan et al., 2024); or learning skinning weights from a set of annotated rigged characters (Xu et al., 2020; Qin et al., 2023). These approaches, however, require various levels of intermediate supervision, are limited in their handling of motion or shape variations, and suffer from limited generalization.
Given a coarse motion specification, we transfer animation to a full-character mesh, without access to any rig or morphable template, at train or test times. A natural approach is to treat this as a sequence prediction problem. However, this quickly leads to memory issues as the animations’ length or the meshes’ resolution increases. Increasing the network capacity adversely affects the situation, leading to overfitting as we often have sparse/limited training data. Also, treating the problem at the frame level is efficient, but leads to jittery motion without any temporal coupling.
Inspired by the recent success of latent space diffusion models (Rombach et al., 2021) over pixel space diffusion models, we ask if a similar latent space can be designed for (sparse) motion sequences. To this end, we propose Kinetic Codes, a temporally-informed lightweight motion autoencoder, that we train over a collection of sparse (humanoid) motion sequences across all types of motion. (Since we only rely on keypoints, instead of body meshes/template, we call the latent space kinetic instead of kinematic.) Regularized by this latent space, we train a spatial and a temporal gradient predictor network. We couple the networks through differentiable spacetime integration and supervise the framework, in an end-to-end fashion.
By representing source motion using only keypoints, we eliminate the need for geometric constraints such as 2-manifoldness, watertightness, or fixed triangulation in source meshes. Moreover, our motion representation allows for 2D source motion as input, which can be transferred to any stylized character (see Figure 1). This keypoint-based representation simplifies motion capture and facilitates sampling and interpolation of motions.
We evaluate SMF for generalization across diverse shapes and unseen motion, and compare against various alternatives. We evaluate our setup on a range of diverse shapes and motion datasets (e.g., AMASS, D4D, Mixamo, monocular video). In summary, we: (i) propose a self-supervised animation transfer framework regularized by kinetic codes, a learned temporally aware latent space; (ii) develop a rig- and template-free animation transfer framework based on simple keypoints as input that is easy to train with sparse supervision and generalizes robustly to new motions and stylized characters; and (iii) report a new SoTA on the AMASS dataset and show realistic motion transfer to in-the-wild stylized characters using different 3D as well as 2D coarse space motion specifications.
Code, weights, and supplementary are available on the project webpage at https://motionfields.github.io/.
2. Related Works
Encoding shape deformation.
Parameterized deformation approaches represent 2D or 3D shapes through a predetermined function of shared parameters and capture deformations as variations of these parameters. A popular example of such models is morphable models (Egger et al., 2019). Such techniques encompass cages, explicit formulations (Ju et al., 2005) or neural approaches (Yifan et al., 2020), blendshapes (Lewis et al., 2014), skinned skeletons (Jacobson et al., 2014), Laplacian eigenfunctions (Rong et al., 2008), etc. Linking these parameters to the shapes’ surface typically necessitates manual annotation of weights, commonly called weight painting, in 3D authoring tools. Alternatively, with access to sufficient supervision data, data-driven approaches can yield realistic neural rigs, as exemplified by Pinocchio (Baran and Popović, 2007), RigNet (Xu et al., 2020), skinning-based motion retargeting (Marsot et al., 2023; Aberman et al., 2020; Zhang et al., 2023), and skeletal articulations with neural blend shapes (Li et al., 2021a). Unsupervised shape and pose disentanglement (Zhou et al., 2020) proposes learning a disentangled latent representation of shape and pose, facilitating motion transfer using shape codes, dependent on registered meshes and maintaining identical connectivity. To plausibly animate these parameterized shapes over time, the parameters should evolve dynamically, weighing the mesh. Although these methodologies require access to body templates and/or rigs, they can still produce jittery results due to loose coupling of individual frame predictions.
Notably, Skeleton-free pose transfer (Liao et al., 2022) aims to alleviate the need for rigs by treating the character pose as a set of independent part deformations. By learning the skinning weights and deformations associated with each module, it can match the source pose using linear blending of skinning weights. While it achieves impressive results in pose transfer, being a per-frame method, it suffers from artifacts and lacks temporal coherence when applied to animation transfer, as observed in our results. Furthermore, its reliance on both source and target shapes being provided in a rest T-pose presents a practical limitation, as such canonical poses are often difficult to define for non-humanoids. This dependency challenges its classification as a truly rig-free method.
Modeling motion as sequence prediction.
Deep recurrent neural networks are capable of modeling time and shape sequences using LSTMs to predict human joints (Fragkiadaki et al., 2015), generate motion in-betweening (Harvey et al., 2020), and to learn a motion field through time (He et al., 2022). These approaches require access to templates/rigs, and large datasets of joint motion since they are discrete time representations. Qiao et al. (2018) utilize mesh convolutions with LSTMs to deform vertices through time; while, Motion Diffusion (Raab et al., 2023) uses local attention to capture motifs of a single motion and combines it with a diffusion UNet module to produce motion extrapolation and in-betweening. The main challenge is handling long (extrapolation) sequences while still being able to generalize to unseen motion. In this work, we propose using Augmented Neural ODEs (Dupont et al., 2019), operating through temporally-aware kinetic codes, to model time continuously instead of using discretized sequential networks such as LSTMs.
Modeling motion using morphable templates.
Temporal surface effects can be modeled, physically correctly, by simulating the underlying soft tissues with finite element methods (Chadwick et al., 1989; Fan et al., 2014). However, this direct simulation is often slow and requires artists to design the underlying bone and muscle structure (Abdrashitov et al., 2021). To overcome the stiffness problem and speed up the simulation, reduced-order models have been proposed (Modi et al., 2020; Park and Hodgins, 2008). When character rigs are available, approaches have been proposed to add soft tissue deformation as an additive per-vertex bump map on top of a primary motion model. Santesteban et al. (2020) use this approach, AMASS (Mahmood et al., 2019) imparts secondary motion using the blending coefficients of the SMPL shape space (Loper et al., 2015), while Dyna (Pons-Moll et al., 2015) learns a data-driven model of soft-tissue deformations using a linear PCA subspace. However, these methods require access to primary motion via a skeleton rig and are restricted to humans registered to a canonical template.
Temporal Residual Jacobian (TRJ) (Muralikrishnan et al., 2024) uses NJF and ODE for motion retargeting, and demonstrates good generalization across shapes. However, TRJ requires motion annotation and also access to SMPL template during training, has to be retrained for different motion classes, produces jitters due to per-frame prediction, and does not generalize to unseen motion. Unlike TRJ, which requires a template to learn motion transfer between different shapes, our self-supervised method reconstructs the original motion on the same shape, eliminating the need for annotated data.
3. Algorithm
Our goal is to animate unrigged triangulated meshes of 3D characters, conditioned on coarse motion signal. These motion control parameters are defined at a few keypoints of the body, can be varied in representation (e.g., 2D or 3D), and specify the target pose per frame. We learn to map these coarse keypoint parameters to dense 3D meshes and, as output, generate an animated 3D mesh at every frame described by the input motion. This nullifies the need for any consistent mesh template or fixed triangulation for both the source motion as well as the target character. Additionally, our method generalizes to unseen motion targets and unseen body shapes, learns from sparse datasets containing a mix of motion examples, and can be applied to long motion sequences.
3.1. Overview
Our method is a module named Self-Supervised Motion Fields () that maps inputs describing a target shape, , and motion, , to per-frame motion as 3D meshes,
where denotes the mesh vertices at frame .
Input representation.
Our inputs are coarse motion parameters, defined as , which characterize the pose required in each frame and the target triangulated mesh defined as . In our experiments, we have tested ranging from 200 (short) to 4000 (long) sequences. We define at fixed keypoints extracted automatically from a given body. Specifically, is a vector at each time step, where is the number of keypoints/joints and is the dimensionality of the chosen input representation. This representation can be 3D keypoint locations on a mesh, 2D keypoints defined on a stick figure frame, or 3D relative Euler angles computed according to the kinematic tree of the chosen pose space. In this work, we focus on results with 2D and 3D keypoints, as we observe their performance to be more robust than 3D relative Euler angles. Our method is self-supervised as we automatically extract the keypoints from mesh sequence.
We now define our module from Equation 1 as,
where describes additional geometric features of the target character such as centroids, normals, and Wave Kernel Signatures (Aubry et al., 2011) computed on the faces of and encoded per-face using a shallow PointNet (Qi et al., 2016). These additional features establish correspondence during inference on unseen in-the-wild shapes. We jointly train the PointNet network and other networks in our system. Figure 2 presents an overview of our method.
Keypoint extraction.
We semi-automatically extract the keypoints . We start with a one-time manual annotation of one human and one animal shape, performed by selecting mesh faces at joint locations. These sparse annotations are then automatically propagated across all characters and motions using point correspondences computed with Diff3F (Dutt et al., 2024). Our framework is
robust to variations between character setups; for example, it bridges the gap even when target keypoint locations (e.g., from Mixamo) differ from the source. This flexibility is possible because our method only requires the number and order of keypoints to be consistent, not their exact spatial positions. This process, where we annotate just two shapes once, allows us to robustly handle diverse character and motion datasets. For extracting 2D keypoints from RGB videos, we use HRNet (Sun et al., 2019), a pretrained pose extractor.
3.2. Kinetic Codes: Temporally-Informed Motion Representation
At the core of our method is the shape deformation module . We observed that naively passing the extracted coarse motion parameters to impairs training, and leads to poor generalization and artifacts on unseen motions (see ablation in Section 4). This is unsurprising as an individual does not contain any temporal motion information or context. Therefore, we first embed in the latent space of an autoencoder. By coupling information across time, this representation leads to smoother interpolation, thereby enhancing generalization to unseen motions.
Further, given the sparsity of training data, we find that using the displacements of the motion parameters as inputs, instead of their absolute values, significantly boosts generalization to unseen motion. Therefore, we express as displacement vectors with respect to the first frame motion parameter as,
We then train a multi-headed attention auto-encoder with self-attention to reconstruct as,
where , are multi-headed attention encoder and decoder networks, respectively; is the per-frame latent motion representation, referred to as kinetic code, of the same dimensionality as , and are the decoded motion displacements; refers to number of frames.
The length of the sequence can vary across motion samples. Note that we use attention on the full sequence as context, and hence a single frame has an understanding of the broader motion sequence.
We train the auto-encoder by minimizing the reconstruction loss,
Thus, we obtain a temporally-informed latent motion representation , which is more conducive to interpolation, resulting in improved generalization to unseen motions and unseen shapes. For our deformation module (described next), we only use as the per-frame motion representation, and freeze and .
3.3. 3D Shape Posing via Deformation Module
Inputs to our shape deformation module comprise the initial configuration, specifically the shape in its canonical rest pose, the learned motion latents , and the geometric features (derived from ). The module is designed to forecast the pose for each time step . While could be trained to directly forecast the target vertex positions, such an approach results in artifacts, including flipped and folded faces. Following neural gradient space processing (Aigerman et al., 2022; Muralikrishnan et al., 2024), we use a simple MLP to predict affine transformation matrices, labeled as Jacobians, at the centroids of the mesh faces, thus encoding relative transforms (Sumner and Popović, 2004). Specifically, we predict final vertex positions using predicted Jacobians, integrated through a differentiable Poisson solver to solve a system of linear equations. We supervise this using vertex-to-vertex and Jacobian losses to ensure accuracy in matching ground-truth data; this frees us from requiring additional annotation data. This improves shape consistency and can model various deformations by predicting affine transformations to mesh faces. Concretely, we define,
where is the Jacobian of the initial frame shape and are intermediate residuals predicted by , which are added to to produce the Jacobian for the frame.
We found that predicting the Jacobians using the Kinetic codes, coupled with residual connections leads to significantly faster and well-behaved convergence and better generalization. Note that is the identity transformation projected on the local basis of each face of by rotation. All Jacobians are predicted in the central coordinate frame defined by the local bases of the faces of .
We can now directly obtain final vertex positions using the Poisson Solver on the predicted Jacobians. However, as we drift away from the pose of the initial shape or when inferring on unseen motion parameters, accumulated error leads to shape inconsistencies and artifacts. Additionally, utilizes only coarse motion parameters and, by itself, is unable to produce temporally coherent and smooth motions on dense meshes for long sequences. Hence, we use an Augmented ODE based formulation, described next, to enable temporally coherent predictions.
3.4. Temporally Coherent Motion Prediction
Motion prediction.
We introduce a second stage that learns to predict a dynamic correction to the initial pose estimate to enforce temporal coherence. We predict the sequences in chunks of consecutive frames, i.e., the given sequence is split into fixed windows, each of size ( in our tests). We set the initial state of the NODE as the Corrective Jacobians required for the first frame as,
To make this operable with ANODE (Dupont et al., 2019), we augment the initial state with extra dimensions. We follow the recommendations of the original work and set the extra dimensions to where denotes the number of extra (augmented) dimensions to be added. We use for all our experiments. In essence, these augmented dimensions act as a form of temporal memory for the ODE. By lifting the state space to a higher dimensional space where it can more easily predict the required trajectories, the system is able to maintain a richer state representation of the motion’s history. This simplifies the task of learning the corrective dynamics needed to prevent drift. In other words, temporal message passing helps predict the corrective Jacobians. Equation 8 becomes,
We predict the correctives in the augmented space first, which is driven by an MLP that predicts the rate of change of the correctives in time. The function models the rate of change of the drift based on conditioning factors,
where , as defined previously, are the Jacobians of the first frame; is the attention encoding of the current window of Jacobian predictions from Equation 7 and is the attention encoding of the previous window of corrective predictions. Intuitively, provides the context of the current motion’s structure, while informs the model about the accumulated error from the previous window. We integrate the local changes over time using Euler’s method to obtain at each time as,
Since is in the augmented space, we use a final linear projection with learnable weights () to map it back to the original unaugmented space. Specifically,
where is simply the learnable weights of the last linear layer projecting from to the Jacobian.
Our jointly trained attention encoders are defined as,
where and are multi-head attention networks, and are the block of sequential Jacobians in the current window and the past window , respectively; and are the corresponding blocks of time instances in these windows, which are positionally encoded.
These attention networks encode a window of Jacobians into a single encoding as shown in Figure 3. Fixing the encoding sizes to a constant size enables handling any arbitrary window/sequence length without overflowing memory. The encoders distill the current window of Jacobian predictions and previously predicted window of Corrective Jacobians . We pass the output of the attention networks as conditioning to Equation 10 to integrate and obtain the correctives in Equation 11 in the augmented space, before projecting them in Equation 12 to obtain the final correctives. We add the predicted correctives to the posed Jacobians from Equation 7. Finally, the predicted Jacobians are spatially integrated using a differentiable Poisson solve (Nicolet et al., 2021), in the coordinate frame of the first frame, to obtain the predicted shape at frame number .
Loss terms.
We train end-to-end using only a shape loss over vertices of and a Jacobian loss. No extra annotation is required for supervision. Our final objective function is,
aggregated together into the total loss as with in our experiments.
4. Evaluation
We test SMF along multiple axes: (i) generalization to unseen motion; (ii) animation transfer to different characters (shapes and topology; meshes and scans); (iii) motion specification using 3D keypoints, using monocular videos, or sampling/interpolating in the Kinetic Code (latent) space. No rigs or templates were used for any training or tests. We also test motion transfer from models trained on humanoid characters to non-humanoids.
| Method | One Leg Jump | Chicken Wings | Walk | Shake Shoulders | Knees | Shake Arms | Shake Hips | |||||||||||||||||||||
| L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | |
| Skeleton-free transfer (Liao et al., 2022) | 4.60 | 0.67 | 0.44 | 13.67 | 3.27 | 0.65 | 0.50 | 13.78 | 5.27 | 0.78 | 0.65 | 19.27 | 5.27 | 0.72 | 0.46 | 13.97 | 4.71 | 0.62 | 0.40 | 12.82 | 3.74 | 0.70 | 0.52 | 15.39 | 3.83 | 0.75 | 0.56 | 16.19 |
| (3D) w/o encoding | 4.56 | 0.70 | 0.32 | 12.56 | 4.22 | 0.68 | 0.35 | 12.86 | 5.84 | 0.71 | 0.34 | 12.92 | 4.19 | 0.68 | 0.30 | 11.85 | 5.20 | 0.77 | 0.36 | 13.79 | 5.21 | 0.80 | 0.44 | 19.79 | 4.48 | 0.71 | 0.35 | 13.26 |
| (2D keypoints) | 5.74 | 0.72 | 0.33 | 12.42 | 6.50 | 0.80 | 0.44 | 19.43 | 6.87 | 0.79 | 0.36 | 13.49 | 6.77 | 0.87 | 0.38 | 15.63 | 8.33 | 0.88 | 0.39 | 14.81 | 8.81 | 0.96 | 0.50 | 22.39 | 7.74 | 0.89 | 0.43 | 17.06 |
| (3D keypoints) | 2.79 | 0.51 | 0.25 | 9.82 | 3.74 | 0.61 | 0.34 | 12.72 | 3.73 | 0.54 | 0.27 | 10.24 | 3.09 | 0.58 | 0.28 | 11.44 | 3.78 | 0.63 | 0.30 | 11.27 | 4.41 | 0.69 | 0.41 | 18.45 | 3.40 | 0.58 | 0.31 | 12.28 |
| Method | Hiphop | Shuffling | Surprised | Shaking Hands | Arguing | |||||||||||||||
| L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | L2-V | L2-V | L2-J | L2-N | |
| Skeleton-free transfer (Liao et al., 2022) | 12.41 | 4.61 | 1.08 | 32.28 | 13.72 | 3.81 | 0.83 | 25.51 | 13.24 | 2.79 | 0.70 | 21.14 | 17.44 | 2.75 | 0.67 | 20.48 | 14.35 | 2.60 | 0.69 | 20.85 |
| (Aigerman et al., 2022) | 25.57 | 9.03 | 1.13 | 45.94 | 25.38 | 9.02 | 1.11 | 45.67 | 26.27 | 8.96 | 1.10 | 45.24 | 31.80 | 10.68 | 1.19 | 45.80 | 29.22 | 9.71 | 1.14 | 45.26 |
| (Muralikrishnan et al., 2024) | 15.23 | 5.86 | 0.87 | 35.09 | 16.91 | 6.62 | 0.85 | 33.54 | 14.22 | 5.20 | 0.64 | 24.95 | 15.99 | 5.94 | 0.73 | 30.25 | 15.17 | 5.39 | 0.70 | 28.62 |
| (3D keypoints) | 8.35 | 3.58 | 0.62 | 24.38 | 7.97 | 3.50 | 0.60 | 23.46 | 4.60 | 2.00 | 0.28 | 10.36 | 5.78 | 2.31 | 0.43 | 17.86 | 5.98 | 2.27 | 0.40 | 16.03 |
| Seen Motion | Unseen Motion | ||||||||||
| Running | Jumping Jacks | Punching | Jiggle on Toes | Shake Hips | Hips | Shake Shoulders | One Leg Jump | Shake Arms | Walk | Chicken Wings | |
| Error (in cm) | 7.73 | 11.88 | 12.03 | 11.13 | 6.88 | 8.36 | 8.34 | 8.51 | 7.95 | 8.96 | 7.89 |
Motion datasets.
We train our method (and baselines) on a single dataset comprising of human motion categories from the AMASS dataset (Mahmood et al., 2019) with each category containing approximately 6-7 motion sequences. We evaluate our method on unseen motion categories from AMASS, sampled motions from Mixamo, animal motion sequences from DeformingThings4D dataset (Li et al., 2021b), and in-the-wild monocular video recordings. The AMASS dataset utilizes the SMPL body model (Loper et al., 2015), which enables generation of motion sequences for diverse body-shapes by varying the shape parameter, . Note that ours does not use this SMPL information during training or inference.
To show that our method works on animals, we train our method on motions from the DeformingThings4D dataset (Li et al., 2021b), which provides animal 4D meshes as deforming sequences and test it on unseen motion sequences.
We train SMF on 9 motion categories from AMASS, each consisting of sequences performed by 6 humans. The number of frames varies from 150-800. For 2D keypoints, we use 3 YouTube videos (2 characters in total), totaling 1hr. Although the videos are noisy as the camera angles change, SMF successfully learns due to the Kinetic codes setup.
4.1. Baselines
We compare our method against recent per frame (pose transfer) methods: NJF (Aigerman et al., 2022) and Skeleton-Free Pose Transfer (Liao et al., 2022). Due to unavailability of pre-processing code for Skeleton-Free Pose Transfer, we use their pretrained models which were trained on substantially more data including AMASS and Mixamo motions as well as stylized characters. Additionally, it requires the full mesh sequence along with T pose meshes for source and target characters. In comparison, ours only takes in sparse keypoints as input. We also compare ours with animation transfer methods: TRJ (Muralikrishnan et al., 2024) (which uses template) and an ablated version of our method without the Kinetic Code () encoding. In the table below, we highlight the differences between SMF and baseline methods.
| Method | Input Requirements |
| NJF (Aigerman et al., 2022) | SMPL Pose Parameters |
| Skeleton Free (Liao et al., 2022) | T-Posed Source & Target characters, full body source motion |
| TRJ (Muralikrishnan et al., 2024) | SMPL Pose and Bodyshape Parameters |
| SMF (ours) | Sparse Keypoints |
| Modified & trained with Keypoints | |
| Modified & trained with Keypoints |
4.2. Metrics and Target Shapes
We evaluate our animation transfer with four main metrics:
Vertex-to-vertex error (L2-V): Measures the Euclidean distance between ground-truth and predicted mesh vertices, indicating how well the global motion is captured.
Velocity error (L2-V): Quantifies differences in vertex velocities across frames, capturing the temporal smoothness and cohesion of the animation.
Jacobian error (L2-J): Assesses deviations in local transformations, revealing unintended deformations.
Angular error of surface normals (L2-N): Calculates the angle between predicted and ground-truth normals, indicating preservation of local surface orientations.
Target Shapes
We evaluate our method on diverse target shapes varied body shapes sampled from SMPL models, human scans from the FAUST dataset (Bogo et al., 2014), characters from the Mixamo library (e.g., skeleton zombie, triceratop, wolf), and in-the-wild meshes from online 3D repositories (e.g., alien, hole-man). As preprocessing, when applicable, we fixed non-manifold meshes.
4.3. Qualitative Results
Generalization to unseen motion and shape.
We present video results on various unseen motion and unseen shapes on our supplementary webpage. Our method produces consistently better generalization to unseen motion categories compared to NJF and TRJ, both of which result in shape distortion and erroneous displacements as they struggle to follow the input (unseen) motion as seen in Figure 4. This is particularly highlighted in scenarios with large displacements, e.g., feet and hands are widened or stretched thin. Note that unlike in the original TRJ (Muralikrishnan et al., 2024), where specialized models were separately trained for each motion type, we retrained a single TRJ, across all the motion types.
SMF also generalizes to new shapes of varied body types, including non-humans despite being trained only on humans. This correspondence from coarse keypoints to a dense mapping across diverse shapes is learned during the self-supervised motion transfer setup and proves to be even capable of generalizing to multi-legged creatures. It preserves the source motion and target shape and the resultant motion is realistic and free from jitters/artifacts. We show comparative results on unseen motion transfer to in-the-wild target meshes in Figure 4. For high-genus shapes, such as the mesh with holes (right half of Figure 4), NJF distorts the shape; TRJ fares comparatively better, it is still riddled with artifacts. Our Kinetic codes not only preserve the target shape but also more faithfully adhere to the motion. See supplemental webpage for video results.
Utilizing a latent representation for motion encoding with smooth interpolation properties leads to improved generalization. Hence, we do not see any significant artifacts even when operating on in-the-wild unseen target shapes for unseen motion. Moreover, we notice without our motion encoding, regions around joints may show melting (see around feet in Figure 6). The same holds true for more accurate bending of the joints (see Figure 4). We note that the AMASS motions contain foot-skating artifacts, and ours faithfully reproduces them. However, we also tested ours on high-quality Mixamo motions as shown in Figure 4 and supplemental webpage, where foot skating is less pronounced. Despite being trained only on AMASS, our model successfully transfers these clean Mixamo motions without introducing any additional skating artifacts. This demonstrates the model’s generalization not only to unseen motions but also to different levels of motion quality. This further suggests that our self-supervised setting is promising and training on higher-quality motion samples would likely yield better results.
Figure 1 and Figure 8 show more animation transfer examples to different humanoids and non-humanoids target meshes, sourced from online in-the-wild meshes and the SHREC’07 dataset (Giorgi et al., 2007). Moreover, Figure 1 showcases support for different types of coarse motion specifications. In Figure 7, we show retargetting results on target animal shapes from the D4D dataset.
Generalization to Mixamo sampled motions on stylized characters.
To transfer motion from Mixamo (Mixamo, 2025), we map its joint locations (3D keypoints) to our system, (note the mapping is not perfect due to misalignment between joint locations, requiring some approximation). Our results demonstrate strong generalization, transferring these motions even to different characters. Notably, TRJ (Muralikrishnan et al., 2024) and NJF (Aigerman et al., 2022) are unable to operate on Mixamo motion, as they strictly depend on the SMPL template, which significantly limits their flexibility. Therefore, we modify their architecture to handle 3D coordinates and remove the shape parameter () module from TRJ. Please see supplemental videos on the webpage for comparison. We further evaluate it on 5 different motions on 3 stylized characters with ground truth generated from Mixamo. As seen in Table 2, we see large improvements over existing methods. Except Skeleton-free transfer (additionally trained on Mixamo data and stylized characters), all other baselines have been trained solely on the AMASS dataset with human characters.
Generalization to monocular capture.
can directly transfer raw motion sequences from 2D input videos captured on handheld cameras to a 3D target shape. We use our formulation based on 2D keypoints to directly operate on 2D input frames. This is a particularly challenging problem as we are lifting motion from 2D to a 3D shape without any explicit supervision or template shape. For our 2D training setup, we additionally train only the autoencoder on in-the-wild videos of exercise videos totaling 1 hour. We note this is noisy but this helps the generalize to RGB captures better as all our 2D renderings of AMASS sequences have a similar camera setup. We extract 2D keypoints using HRNet (Sun et al., 2019), which computes 2D keypoints frame by frame. We use Savitzky-Golay filter to smoothen the 2D signal as there is inter-frame noise.
Self-supervised training.
Our self-supervised setup enables us to learn motion transfer without requiring a template shape and using only coarse motion parameters extracted at joints from the motion dataset. This setup is the key to enabling learning from coarse input motion (2D/3D keypoints) and transferring motion from 2D monocular captures and Mixamo sampled motion to 3D target shapes.
4.4. Quantitative Comparison
We present quantitative results on unseen motion category and unseen motion sequence (within the same category) to unseen shape in Tables 1 and 2. Generalization to unseen motion is a particularly challenging problem as the Poisson solve in NJF and TRJ facilitates shape preservation for unseen shape by following the coordinates of the source shape. However, there is no such aid when generalizing to unseen motion, which is an extrapolation problem. Our method significantly outperforms NJF and TRJ for both motion accuracy as well as detail preservation. The benefits of our are highlighted here as we see significantly lower errors against a baseline of without encoding. Compared to the per-frame pose transfer performed by Skeleton-Free Pose Transfer which has numerous artifacts and unnatural deformations, SMF also ensures temporal consistency as seen in lower velocity error (L2-V), measuring motion smoothness.
Error accumulation on long motion transfer.
We choose 3 unseen, complex, and long dance sequences (10000 frames) to analyze error accumulation. We report per frame vertex-to-vertex (L2-V) error in Figure 9. Compared to the only other temporal method, TRJ, SMF does not exhibit a trend of monotonically increasing error demonstrating its robustness to temporal drift and its suitability for generating long motion sequences. The initial spike in error is due to the motion moving away from the easy-to-model rest pose.
4.5. Performance
SMF’s lightweight architecture enables inference at ¿30 FPS on a single RTX 4090 for a mesh with 30k faces (for high number of faces this will be slower). This makes our system suitable for real-time applications, following a one-time preprocessing step to compute features on the target mesh. Furthermore, when using 2D keypoints, the system could potentially be extended for real-time 2D-to-3D animation transfer, as keypoint extraction with HRNet (Sun et al., 2019) is highly efficient.
5. Conclusion
We have presented Self-supervised Motion Fields that convert sparse motion into full-body character motion. We enable this by first creating a temporally-informed motion latent space – Kinetic codes – and then utilizing it to train spatial and temporal gradient predictor networks jointly. The gradient signals are coupled via spatial and temporal integration, and trained using full-body mesh sequences for supervision. SMF does not require additional motion annotation, is simple and robust to train, and generalizes across unseen motion and character shape variations. We also do not use any template model or rig, either during training or inference. We establish a new SOTA on the AMASS dataset.
Limitations. While our formulation captures some secondary motion, we do not model secondary dynamics in any physically correct way. Neither do we explicitly support collision handling to prevent self-intersection (see the triceratop leg crossing in Figure 7). An interesting future direction would be to implicitly handle collision detection and simultaneously model character-garment interactions, possibly using a transformer-based attention mechanism to capture non-local interactions between body parts. Similar to single human bodies (Frühstück et al., 2022), it will be interesting to combine different motion generators for different parts of a character, allowing mixing of motion modes. Finally, temporal sketching sequences (Fu et al., 2011) can also be modeled with such a generative framework.